OpenStudy (maeganr):

PLEASE HELP! : Question: "Walnuts cost $6 per pound and macadamia nuts cost $2 per pound. You have $20 to spend on nuts for baking. What are three possible combinations of walnuts and macadamia nuts you can buy?" PLEASE help me show my work and work through it!

OpenStudy (maeganr):

PLEASE HELP ASAP!

OpenStudy (maeganr):

I will give out medals(:

OpenStudy (anonymous):

lets make some variables... w = walnuts m = macademia so lets make an equation based on those variable and your given information...

OpenStudy (maeganr):

$6w+$2m=$20 ?

OpenStudy (anonymous):

good..

OpenStudy (maeganr):

Now what do I do?

OpenStudy (anonymous):

without more information...now must use logic how many w and m can we get to total 20 based on that equation?

OpenStudy (maeganr):

But dont you have to show work and do it solving the equation?

OpenStudy (anonymous):

for instance, if w =3 than she is able to have 1 m 6(3) + 2(1) = 20 you try one

OpenStudy (maeganr):

6(2)+2(4)=20

OpenStudy (anonymous):

did they provide more information, based on what you typed...can't create a 2nd equation

OpenStudy (anonymous):

good one...

OpenStudy (maeganr):

Just what I typed is all they gave me.

OpenStudy (anonymous):

then the way we are doing is the only way....

OpenStudy (maeganr):

Ok, so whats a 3rd combination?

OpenStudy (anonymous):

6(1) + 2(7) = 20

OpenStudy (maeganr):

Thank you! I will post another question soon..Can you help? Ill tag you..